Determine the nature of the solutions of the equation.
y^2 = 3/2y + 2/7
a. 1 real solution
b. 2 non-real solutions
c. 2 real solutions
I am confused about the fractions number one and how to determine how many real or non-real solutions there are. Can someone please help?
This is not from a textbook.

Answer by stanbon(75887)

y^2 = 3/2y + 2/7
Rearrange:
y^2-(3/2)y - (2/7)
a= 1, b = -3/2 , c = 2/7
discriminant: b^2 - 4ac
(3/2)^2 -4*1*(-2/7)
9/4 + 8/7
(63+32)/(4*7)
= 95/28
This is greater than zero so there are
2 real solutions.
-------------------------
Cheers,
Stan H.
=========================
a. 1 real solution
b. 2 non-real solutions
c. 2 real solutions